The IS and LM curves for the economy have the following equations: Y = k (Ap - 50 i)

Y = 2(Ms / P) + 100 i

where k = 2, Ap = 4000, Ms = 3000 and P =1

Find the equilibrium level of income and the equilibrium interest rate. Complete following table which represents points on aggregate demand curve.

Price Interest Rate Real GDP

.8 __________

1 _______________________

1.5 ____________

2.0 _____________
Answer #1

Page 1 O f IS: Y = K (Ap- soi) LM: Y: 2 M P + loo e putting the values for R, Ap, MS and taking - 1 at equilibreuim, IS = LMPage 2 & IS: Y = Cap-50 :) LM: Y 22 (Ms / P) + 100%. ③ With P=0.8. of equubelium IS-LM K(Ap-soi) = 2(Ms /P) + 100. 2 (84000lage-31 2C4000-501):9(8000/15) 100 9000 - 1001 - 4000 41001 2ooi = 4000 i - 40001 - 20 i -20 putting value of i in een o y=9[Page 4 -i-25 putting value of I in een 0. Y = 2 [4000 - 5o (25)] 3a 14000 - 1250) 2 2x 2750 2 5500 1 Y=5500 Completing the ta